# Multiscale interpolative construction of quantized tensor trains

### v2

@@ 1 Introduction
However, this type of analysis cannot fully explain the approximation power of QTTs. First, it cannot explain why the ranks tend to decay with increasing depth in the QTT, since rank bounds derived from summation of complex exponentials or polynomial approximation are uniform across all depths. Second, it cannot explain why certain functions with sharp features (i.e., poor quantitative smoothness) often admit low-rank representations as QTT. We comment that some recent work [2, 1] has studied quantized tensor networks through the lens of Besov spaces, though the goals and focus of that work are quite different from our own.
In this work, we analyze the construction of QTTs from the point of view of multiscale polynomial interpolation. Our analysis addresses both of the open questions highlighted above.
First, we show quantitative rank bounds in terms of the smoothness of the target function which decay with depth. In fact, our decaying rank bounds imply that $\Omega$-bandlimited functions have QTT ranks uniformly bounded by $O(\sqrt{\Omega})$. (See Corollary 10 below for a precise statement.) This result is particularly striking because, by expressing an $\Omega$-bandlimited function as a linear combination of $O(\Omega)$ complex exponentials, we naively expect a uniform rank bound of only $O(\Omega)$.
Second, we reach the stylized conclusion that functions which are well approximated in a multiresolution polynomial basis are well represented as QTTs.
@@ 1.1 Outline
In Section 2, we review the idea of QTTs. In Section 3, we prove rank bounds for QTTs that decay with depth. In Section 4, we present our practical approach for QTT construction, first in its most basic form but then with several extensions that improve the practical efficiency. In Section 5, we explain how to ‘invert’ the QTT construction, recovering function evaluations on interpolating grids from a given QTT. In Section 6, we present a multiresolution extension of our interpolative construction, as well as theory that explains the representability of certain functions with sharp features as QTT. In Section 7, we explain how the preceding discussion extends to the setting of multivariate functions, where several conventions for ordering tensor indices are possible. In Section 8, we conclude with numerical experiments illustrating our theory and practical algorithms.
@@ 3.2 Interpolation bounds
By combining Proposition 8 with the trivial bound for the leading ranks of an MPS/TT, we obtain the striking result that the QTT ranks of an $\Omega$-bandlimited function are uniformly bounded by $O\left(\sqrt{\Omega}+\log_{+}\left(\frac{|\mu|}{\varepsilon}\right)\right)$.
@@
Let $T$ be a quantized tensor representation for $f$ on $\mathcal{D}_{K}$, where the depth $K$ is an arbitrary positive integer and $f:[0,1]\rightarrow\mathbb{R}$ is the restriction of an $\Omega$-bandlimited function with spectral measure $\mu$. Assume without loss of generality that $\Omega\geq 1$. Then for any $\varepsilon>0$, the $(\varepsilon,\infty)$ ranks of all unfolding matrices of $T$ are bounded uniformly by $O\left(\sqrt{\Omega}+\log_{+}\left(\frac{|\mu|}{\varepsilon}\right)\right)$.
@@
Consider the limit of large $\Omega$. Since the Shannon sampling theorem suggests that we can resolve $f$ up to fixed accuracy with a trigonometric polynomial of order $O(\Omega)$, by constructing $f$ as a sum of $O(\Omega)$ complex exponentials, we might naively expect that the QTT ranks of $f$ are $O(\Omega)$. However, Corollary 10 in fact suggests that the ranks are $O(\sqrt{\Omega})$.
@@
$r_{m}^{(\varepsilon,\infty)}[T]\leq 1+\left\lceil\sqrt{\Omega}+2\log_{+}\left(\frac{2|\mu|}{\pi\varepsilon}\right)\right\rceil=O\left(\sqrt{\Omega}+\log_{+}\left(\frac{|\mu|}{\varepsilon}\right)\right)$
for $\Omega\geq 1$. Therefore we have the desired rank bound for all $m$. ∎
@@ 4 Interpolative construction of QTTs
Although we have shown how smoothness quantitatively bounds the ranks of the unfolding matrices of a quantized tensor representation, we have not demonstrated an explicit construction of a QTT. In Section 4.1, we will present a direct construction based on Chebyshev interpolation using the Chebyshev-Lobatto grid $c_{N}^{\alpha}=c^{\alpha}$, where we omit $N$ from this notation going forward for visual clarity. The TT ranks in this construction are all $N+1$.
@@
Finally, in Section 4.3, we show how the dense tensor cores implementing Chebyshev interpolation can be replaced with sparse tensors, following [3], which further reduces the computational cost to $O(Nr^{2})$.

### v3

@@ 1 Introduction
However, this type of analysis cannot fully explain the approximation power of QTTs. First, it cannot explain why the ranks tend to decay with increasing depth in the QTT, since rank bounds derived from summation of complex exponentials or polynomial approximation are uniform across all depths. Relatedly, we will see that even the global rank bounds derived from such approaches are suboptimal. Second, it cannot explain why certain functions with sharp features (i.e., poor quantitative smoothness) often admit low-rank representations as QTT. We comment that some recent work [2, 1] has studied quantized tensor networks through the lens of Besov spaces, though the goals and focus of that work are quite different from our own.
In this work, we analyze the construction of QTTs from the point of view of multiscale polynomial interpolation. Our analysis addresses the open questions highlighted above.
First, we show quantitative rank bounds in terms of the smoothness of the target function which decay with depth. In fact, our decaying rank bounds imply that $\Omega$-bandlimited functions have QTT ranks uniformly bounded by about $\sqrt{\Omega}$. (See Corollary 10 below for a precise statement.) This result is particularly striking because, by expressing an $\Omega$-bandlimited function as a linear combination of $O(\Omega)$ complex exponentials, we naively expect a uniform rank bound of only $O(\Omega)$.
Second, we reach the stylized conclusion that functions which are well approximated in a multiresolution polynomial basis are well represented as QTTs. This conclusion is formalized in Theorem 20.
@@ 1.1 Outline
In Section 2, we review the idea of QTTs. In Section 3, we prove rank bounds for QTTs that decay with depth. In Section 4, we present our practical approach for QTT construction, first in its most basic form but then with several extensions that improve the practical efficiency. (See the beginning of Section 4 for a detailed outline of each subsection.) In Section 5, we explain how to ‘invert’ the QTT construction, recovering function evaluations on interpolating grids from a given QTT. In Section 6, we present a multiresolution extension of our interpolative construction, as well as theory that explains the representability of certain functions with sharp features as QTT. In Section 7, we explain how the preceding discussion extends to the setting of multivariate functions, where several conventions for ordering tensor indices are possible. In Section 8, we conclude with numerical experiments illustrating our theory and practical algorithms.
@@ 3.2 Interpolation bounds
By combining Proposition 8 with the trivial bound for the leading ranks of an MPS/TT, we obtain the striking result that the QTT ranks of an $\Omega$-bandlimited function are uniformly bounded by $\sqrt{\Omega}+O\left(1+\log_{+}\left(\frac{|\mu|}{\varepsilon}\right)\right)$.
@@
Let $T$ be a quantized tensor representation for $f$ on $\mathcal{D}_{K}$, where the depth $K$ is an arbitrary positive integer and $f:[0,1]\rightarrow\mathbb{R}$ is the restriction of an $\Omega$-bandlimited function with spectral measure $\mu$. Then for any $\varepsilon>0$, the $(\varepsilon,\infty)$ ranks of all unfolding matrices of $T$ are bounded uniformly by $\sqrt{\Omega}+O\left(1+\log_{+}\left(\frac{|\mu|}{\varepsilon}\right)\right)$.
@@
Consider the limit of large $\Omega$. Since the Shannon sampling theorem suggests that we can resolve $f$ up to fixed accuracy with a trigonometric polynomial of order $O(\Omega)$, by constructing $f$ as a sum of $O(\Omega)$ complex exponentials, we might naively expect that the QTT ranks of $f$ are $O(\Omega)$. However, Corollary 10 in fact suggests that the ranks are about $\sqrt{\Omega}$.
@@
$r_{m}^{(\varepsilon,\infty)}[T]\leq 1+\left\lceil\sqrt{\Omega}+2\log_{+}\left(\frac{2|\mu|}{\pi\varepsilon}\right)\right\rceil=\sqrt{\Omega}+O\left(1+\log_{+}\left(\frac{|\mu|}{\varepsilon}\right)\right).$
Therefore we have the desired rank bound for all $m$. ∎
@@ 4 Interpolative construction of QTTs
Although we have shown how smoothness quantitatively bounds the ranks of the unfolding matrices of a quantized tensor representation, we have not demonstrated an explicit construction of a QTT. In Section 4.1, we will present a direct construction based on Chebyshev interpolation using the Chebyshev-Lobatto grid $c_{N}^{\alpha}=c^{\alpha}$, where we typically omit $N$ from this notation going forward for visual clarity. The TT ranks in this construction are all $N+1$.
@@
Then in Section 4.3, we show how the dense tensor cores implementing Chebyshev interpolation can be replaced with sparse tensors, following [3], which further reduces the computational cost to $O(Nr^{2})$.
Finally, in Section 4.4, we show how decaying ranks can be built directly into the QTT construction in the special case of bandlimited functions. This section yields an a priori, rather than a posteriori (cf. Section 4.2), guarantee of decaying QTT ranks, nearly matching our corresponding guarantees for the unfolding matrices.
@@ 4.1 Basic construction
Figure 4.1: The core $A$ interpolates values from the Chebyshev-Lobatto grid $\{x_{\leq m}+2^{-m}c^{\alpha}\}_{\alpha=0}^{N}$ for the dyadic subinterval $[x_{\leq m},x_{\leq m}+2^{-m}]$, depicted on top, to the Chebyshev-Lobatto grids $\{x_{\leq m}+2^{-(m+1)}\sigma+2^{-(m+1)}c^{\beta}\}_{\beta=0}^{N}$ for the left $(\sigma=0$) and right $(\sigma=1$) halves of this dyadic subinterval. The interpolation is indicated graphically in the figure for the case $\sigma=0$.
@@
The interpretation of the interpolating tensor core $A$ is illustrated in Figure 4.1.
@@ 4.3 Sparse interpolative construction
4.4 A priori decaying-rank construction
Recall that Proposition 13 guarantees that our basic construction of Section 4.1 is accurate. Note that in this construction, all of the QTT ranks are of the same size (i.e., $N+1$). Meanwhile, Theorem 14 guarantees that truncating the ranks ‘on-the-fly’ to specified tolerance is safe, in the sense that the error of the full QTT is controlled by this tolerance. However, this a posteriori bound does not directly imply an a priori guarantee that the decaying ranks of the unfolding matrices, proved in Section 3, will be revealed.
Ideally, we would like an a priori construction with decaying ranks, matching the rank guarantees for the unfolding matrices. For simplicity we will focus on the bandlimited case, where we have an intuition that the ideal ranks should decrease by roughly a factor of $2$ at each level of the QTT. Therefore we should consider interpolating grids that decay in size by roughly a factor of $2$ at each level.
Concretely, consider $f$ which is $\Omega$-bandlimited in the sense of Definition 7. Then define grid sizes
$N_{k}=\left\lceil 2^{-k}\Omega+\Delta\right\rceil,\quad k=1,\ldots,K,$
where $\Delta>0$ is a free parameter, the choice of which we will discuss later.
Then we define interpolating cores $A_{k}\in\mathbb{R}^{2\times N_{k-1}\times N_{k}}$ for $k=2,\ldots,K$:
$A_{k}^{\alpha\beta}(\sigma):=P_{N_{k-1}}^{\alpha}\left(\frac{\sigma+c_{N_{k}}^{\beta}}{2}\right),\quad\alpha\in[N_{k-1}+1],\ \beta\in[N_{k}+1],\ \sigma\in\{0,1\}.$ (4.11)
Following the shorthand adopted in preceding sections, our QTT will be defined in terms of its cores as
$S=A_{\mathrm{L}}A_{2},\ldots,A_{K-1}A_{\mathrm{R}},$ (4.12)
where $A_{\mathrm{L}}$ is defined as in (4.1) with $N=N_{1}$ and $A_{\mathrm{R}}$ is defined by $A_{\mathrm{R}}^{\alpha}(\sigma):=A_{K}^{\alpha 0}(\sigma)$.
Theorem 16.
Let $S$ denote the MPS (4.12) that is furnished by the above construction, and let $T$ denote the true quantized tensor representation of the target function $f:[0,1]\rightarrow\mathbb{R}$, which we assume to be the restriction of an $\Omega$-bandlimited function with spectral measure $\mu$. The construction error is bounded as
$\|T-S\|_{\infty}\leq\frac{2}{\pi}|\mu|(K-1)\left[1+\frac{2}{\pi}\log(\lceil\Omega/2+\Delta\rceil+1)\right]^{K-2}e^{-\Delta/2}.$
Remark 17.
The factor in the bound with the $K$-dependent exponent results from repeated error amplification by the Lebesgue constant of the interpolating cores. In practice, this aspect of the bound seems not to be sharp, though it appears to be difficult to rule out such a dependence. Nonetheless, this factor can be suppressed efficiently in the bound by only a modest increase of $\Delta$ with mild dependence on the depth $K$.
Remark 18.
Note, following the same reasoning as in the proof of Corollary 10, that the largest QTT rank of this construction is bounded by $\lceil\sqrt{\Omega}+\Delta\rceil+1$, since the leading ranks $k=1,\ldots,\lfloor\frac{1}{2}\log_{2}\Omega\rfloor$ are trivially bounded by $\sqrt{\Omega}$ (the number of rows in the unfolding matrix) and the ranks $k\geq\lceil\frac{1}{2}\log_{2}\Omega\rceil$ are bounded by $\lceil\sqrt{\Omega}+\Delta\rceil+1$.
Proof.
As in Section 4.2, let $S_{\leq m}=A_{\mathrm{L}}A_{2}\cdots A_{m}$ denote our QTT at the $m$-th stage of construction. Let
$\varepsilon_{m}:=\|S_{\leq m}-T_{\leq m}\|_{\infty},$
where $T_{\leq m}$ is defined by
$T_{\leq m}^{\alpha}(\sigma_{1:m})=f\left(\sum_{k=1}^{m}2^{-k}\sigma_{k}+2^{-m}c_{N_{m}}^{\alpha}\right).$
Note that then in particular $\|T-S\|_{\infty}\leq\varepsilon_{K}$, so we seek a bound on $\varepsilon_{K}$, which we build inductively.
Compute:
$\varepsilon_{m+1}$ $=\|S_{\leq m}A_{m+1}-T_{\leq m+1}\|_{\infty}$
$\leq\|(S_{\leq m}-T_{\leq m})A_{m+1}\|_{\infty}+\|T_{\leq m}A_{m+1}-T_{\leq m+1}\|_{\infty}$
$\leq\Lambda_{N_{m}}\underbrace{\|S_{\leq m}-T_{\leq m}\|_{\infty}}_{=\,\varepsilon_{m}}+E_{m,N_{m}}[f],$
where as in the proof of Theorem 14, $\Lambda_{N_{m}}$ denotes the Lebesgue constant for $(N_{m}+1)$-point Chebyshev-Lobatto interpolation. To proceed from the penultimate line to the last line, we used (1) the fact that the Lebesgue constant bounds the $\infty$-operator norm of polynomial interpolation and (2) the fact that the error $T_{\leq m}A_{m}-T_{\leq m+1}$ is a Chebyshev interpolation error precisely in the sense of (3.4).
It follows then, via Proposition 8, that
$\varepsilon_{m+1}\leq\Lambda_{N_{m}}\varepsilon_{m}+\frac{2|\mu|}{\pi}e^{\frac{1}{2}(2^{-m}\Omega-N_{m})}.$
Since $N_{m}\geq 2^{-m}\Omega+\Delta$ by definition, we have
$\varepsilon_{m+1}\leq\Lambda_{N_{m}}\varepsilon_{m}+\frac{2|\mu|}{\pi}e^{-\frac{\Delta}{2}}.$
Then recalling from [22] that
$\Lambda_{M}\leq 1+\frac{2}{\pi}\log(M+1)$
for all $M$, we have
$\varepsilon_{m+1}\leq a\varepsilon_{m}+b,$ (4.13)
where $a:=1+\frac{2}{\pi}\log(\lceil\Omega/2+\Delta\rceil+1)$ and $b:=\frac{2|\mu|}{\pi}e^{-\Delta/2}$.
From (4.13), together with the base case $\varepsilon_{1}=0$, it follows that
$\varepsilon_{K}\leq b\sum_{k=0}^{K-2}a^{k}\leq b(K-1)a^{K-2},$
i.e.,
$\varepsilon_{K}\leq\frac{2}{\pi}|\mu|(K-1)\left[1+\frac{2}{\pi}\log(\lceil\Omega/2+\Delta\rceil+1)\right]^{K-2}e^{-\Delta/2}.$
Since $\|T-S\|_{\infty}\leq\varepsilon_{K}$, this implies the desired bound on $\|T-S\|_{\infty}$. ∎